# A New Causal Rule Learning Approach to Interpretable Estimation of Heterogeneous Treatment Effect

### v2

@@ Abstract
Interpretability plays a critical role in the application of statistical learning for estimating heterogeneous treatment effects (HTE) for complex diseases. In this study, we leverage a rule-based workflow, namely causal rule learning (CRL) to estimate and enhance our understanding of HTE for atrial septal defect, addressing an overlooked question in previous literature: what if an individual simultaneously belongs to multiple groups with different average treatment effects? The CRL process consists of three steps: rule discovery, which generates a set of causal rules with corresponding subgroup average treatment effects; rule selection, which identifies a subset of these rules to deconstruct individual-level treatment effects as a linear combination of subgroup-level effects; and rule analysis, which outlines a detailed procedure for further analyzing each selected rule from multiple perspectives to identify the most promising rules for validation. Extensive simulation studies and real-world data analysis demonstrate that CRL outperforms other methods in providing interpretable estimates of HTE, especially when dealing with complex ground truth and sufficient sample sizes.
@@ 1 Introduction
In clinical research and practice, patients respond differently to certain treatment interventions, necessitating the exploration of heterogeneous treatment effects (HTE) to inform personalized treatment plans and optimize health outcomes. From a practitioner’s perspective, this exploration requires not only accurate estimation but also a deep understanding of HTE, which can be challenging to achieve in real clinical treatment [12], especially for complex diseases such as atrial septal defect (ASD). As a common type of congenital heart disease, ASD presents various pathologies and heterogeneity in patients [8]. With accurate HTE estimation and enhanced interpretation of HTE estimates, cardiologists are able to understand, trust, and justify their treatment recommendations, fostering doctor-patient communication and collaboration [20, 12].
@@
So far, researchers have made great efforts to estimate HTE accurately using various methods, among which statistical learning methods are particularly popular due to their ability to handle high-dimensional datasets and uncover non-linear relationships between covariates [45, 50, 37, 51]. However, most methods are typically designed for general estimation purposes, lacking specific adaptations for clinical applications. Stepping back, they seldom provide guidance on how to interpret what their model has learned from the data. Consequently, clinicians may find it hard to understand these methods, let alone the application of them in real practice.
@@
There has been increasing research to deal with the above interpretability issue in HTE estimation [13, 17, 22, 47, 10, 15], and the majority focuses on identifying subgroups with enhanced (positive) treatment effect. However, the setup and methodology of these studies may be far from sufficient to reveal the complex nature of certain diseases. For complex disease treatment whose outcome is influenced by a lot of factors and interactions among these factors, we may find the situation shown in Figure 1: There are $M$ true subgroups (from subgroup 1 to subgroup $M$) that contribute to treatment effect heterogeneity. The population within each subgroup (e.g., subgroup 1) shares the same treatment effect ($\tau_{1}=20$), i.e., subgroup conditional average treatment effect (CATE) derived from the difference between two treatment groups within the subgroup. The subgroup CATEs between groups vary in effect size (or magnitude) and sign ($\tau_{2}=15$ and $\tau_{3}=-8$). Besides, these subgroups are characterized with (not necessarily) different features, such as subgroup 1 by $chdhis\leq 1$ and $diabp\leq 58$ or subgroup 2 by $NYHA\leq 1$ and $BMI\leq 16.02$. Therefore, when we find the individual belongs to subgroups 1, 2, and $M$ ($r_{1}=1,r_{2}=1,r_{M}=1$) at the same time (multiple memberships), it raises the question of what the individual level treatment effect (ITE) should be, given the three subgroups differ in treatment effect. Considering the complex nature of certain diseases, we assume ITE as a linear combination of the above candidate group CATE in this study. Hence, the ITE in Figure 1 is $1/3\times 20+1/5\times 15+1/10\times(-7)=173/21$ where $\beta_{1}=1/3$, $\beta_{2}=1/5$, and $\beta_{M}=1/10$ are the linear coefficients in the combination which quantifies the contribution of each candidate subgroup to ITE.
@@
Compared with our perspective, the studies focus on subgroup identification usually end up with a limited number of subgroups for interpretability concern [11, 44] (e.g., only subgroup 1 and 2 in Figure 1 are identified). Some approaches even define mutually disjoint subgroups (e.g., [47]): One individual can only belong to one subgroup. As indicated in [44], individual effects are not guaranteed to have even the same sign as CATE, not to mention the same magnitude. Therefore, an ITE cannot be determined solely by belonging to a few specific groups, which relates to a few covariates. On the contrary, ITE may involve a broader range of subgroup memberships, and we need to consider how being in these subgroups collectively influences an individual, i.e., to explore the relationship between individual-level treatment effects and group-level treatment effects. In addition, since these studies emphasize a few subgroups, they are likely to ignore the estimation and understanding of ITE for the population outside the identified subgroups.
@@
With consideration of the above, this study introduces a novel method, Causal Rule Learning (CRL), to integrate both accurate estimation and an in-depth understanding of HTE for complex diseases. Guided by the Predictive, Descriptive, Relevant (PDR) framework of interpretability [31] (see Appendix A for more information) and inspired by the RuleFit [16] methodology, CRL caters to healthcare practitioners via an easily understood rule-based approach and an enlightening perspective on the relationship between ITE and CATE. Specifically, CRL incorporates three steps, i.e., rule discovery, rule selection, and rule analysis. The rule discovery step first generates a set of causal rules (characterized by different patient covariates) and their corresponding subgroup CATE estimates using causal forest [45]. Then, the rule selection step leverages a D-learning method [34] to filter out non-informative rules and use the informative ones to estimate ITE. Lastly, the rule analysis step evaluates the rules from three different granularity levels, namely overall analysis, significance analysis, and decomposition analysis. The whole workflow with details is shown in Figure 2.
@@
Figure 1: Connection between individual level and subgroup level treatment effects.
@@
CRL contributes to the existing literature and healthcare practice by enriching the toolset for interpretable HTE estimation, particularly in scenarios involving complex disease treatments. In detail, we propose:
@@
    A comprehensively interpretable framework for HTE estimation. CRL presents an end-to-end rule-based workflow that includes the discovery, selection, and analysis of rules. Rules are used for their easiness in human understanding, usefulness in healthcare practice, and inclusion of potential knowledge. The discovery step seeks large candidates of rules, which is meaningful for revealing the unknown aspects of HTE. The following selection step imposes sparsity on these rules, surfacing the most predictive and promising ones for ITE and further analysis. The analysis step provides an in-depth investigation of the rules identified, ranging from the overall performance of the rules as a whole on HTE-related tasks to how each component contributes to a specific rule on differentiating effect magnitudes, providing in-depth interpretation of the subgroups identified.
@@
Together, the refined rules, their corresponding subgroup CATEs, and weights (coefficients in the linear combination) obtained after the whole workflow provide insights to clinicians as clinical hypotheses for future validation through clinical trials and large-scale studies, enabling actionable suggestions in healthcare practice.
The rest of this paper is organized as follows: In Section 2, we briefly summarize previous studies on HTE for clinical purposes. Section 3 delineates the proposed CRL workflow, including its basic data settings and assumptions, formulation, and workflow. Section 4 demonstrates the efficacy of CRL on both simulated data and real-world data, respectively. Section 5 concludes our work and discusses its limitations and potential future research directions. For brevity in this paper, we will slightly misuse the following terms: i) HTE, CATE, and ITE; ii) satisfy a rule and belongs to a subgroup. These terms are used according to specific contexts and sometimes indiscriminately in broader situations.
@@ 2 Related work
This section presents extant literature related to HTE estimation from various perspectives. The majority of studies focus on the estimation of CATE, namely the average treatment effect on specific subgroups. A small strand of research aims at treatment recommendation, i.e., directly modeling the best individualized treatment rules or policy given personal characteristics to maximize the outcome of interest. Weighting methods are frequently used for this goal, such as [50, 19, 51, 29]. Though the latter is capable of providing actionable suggestions to clinical treatment decisions, they often fail to give explicit decision rationale or insights into HTE, which is really crucial for complex disease treatment. Therefore, we mainly discuss the studies on CATE estimation below.
CATE research generally has two focal points. One is developing general CATE estimators with favorable statistical properties using statistical methods such as single-index models [38, 14] or (causal) machine learning methods such as causal forest [5] and meta-learners [18, 32, 42]. Some studies involve complicated computations or even black-box models [26, 48, 1] that are inherently unexplainable to improve estimation accuracy, failing to provide sufficient interpretation catering to clinical practitioners. The other focus emphasizes identifying and characterizing subgroups with statistically significant and practically meaningful heterogeneity in treatment response, such as [9, 39]. Tree-based methods such as [17, 30, 37, 39] have a natural advantage for this end due to their heuristic mechanism on population segmentation.
In particular, there is a growing field of CATE studies with rule ensemble models that follow quite different philosophies and workflows. As examples, [13] proposes a method called Stable Discovery of Interpretable Subgroups via Calibration to discover interpretable and stable rule sets from well-calibrated samples across 18 popular CATE estimators. Similarly, [28] uses multiple CATE estimators as outcomes to generate tree ensembles for rule generation and then prune the rules with certain penalties. They re-estimate CATE for the subpopulation characterized by the remaining rules. [47] proposes Causal Rule Sets to identify subgroups with enhanced treatment effect through a generative Bayesian framework that assumes a prior for the actual rule set and a Bayesian logistic regression model to improve it. As mentioned earlier, all these studies presume a limited number of true subgroups which may lead the truth of complex disease under-explored, especially for the population outside the identified subgroups.
@@ 3.1.1 Notation and assumptions
Suppose that we have observed a data set $\mathcal{D}=\{X_{i},A_{i},Y_{i}\}^{N}_{i=1}$ which comprises $N$ independent and identically distributed samples. This data set may come from an observational study or, more ideally, a randomized controlled trial (RCT) setting. The proposed framework can be applied to data from both settings but requires the satisfaction of the three basic assumptions of the Neyman–Rubin potential outcome framework [36, 35], namely the Stable Unit Treatment Value Assumption, Unconfoundedness Assumption, and Overlap Assumption. Details of these assumptions are given in Appendix B.
3.1.2 Individualized treatment effect based on causal rules (subgroup CATE)
In this study, we estimate ITE directly through a set of causal rules. By causal rules, we mean that the rules are identified from a causal analysis framework. As a natural way to represent human knowledge and a widely-used tool in clinical guidelines, they uncover variables’ potential interaction (synergistic effect) and define certain subgroups to help us understand the heterogeneity of the treatment effect. Most importantly, [16] and [31] show evidence that rule-based models have comparable and sometimes even better performance than black box methods like neural networks. Therefore, rules match our needs for both accuracy and interpretability. This paper formulates a rule as:
@@ 3.1.2 Individualized treatment effect based on causal rules (subgroup CATE)
where $x_{j}$ denotes the $j$-th element of $X$, $s_{j}$ is a specified subset of all values of $x_{j}$, and $\mathop{1}(\cdot)$ is the indicator function yielding 1 when $x_{j}\in s_{j}$ and 0 otherwise. For example, $x_{1}>0$ and $x_{2}<0$ and $x_{3}>5$ is a rule representation consisting of three components. Each component (e.g., $x_{1}>0$) has a splitting variable ($x_{1}$) and a given cut-off value or threshold (0), which defines $s_{1}$. A rule $r(X)$ maps an input $X$ into $\{0,1\}$, representing the satisfaction of a rule with 1 and 0 otherwise. Note that a rule is usually represented only by the variables that help to split the whole population, and the ones that do not are often ignored.
@@
where $r_{m}(X)$, $\tau_{m}$, and $\beta_{m}$ denote the satisfaction of the $m$-th rule (0 or 1) given $X$, its corresponding subgroup CATE and weight in the linear combination respectively. Each subgroup CATE has a different contribution to ITE, as reflected by the sign and magnitude of the weights $\mathrm{\beta}$ and the subgroup CATEs. More significant absolute values of these quantities imply more contribution. Given above, the problem then boils down to the revelation of $\{r_{m}(X),\tau_{m},\beta_{m}\}^{M}_{m=1}$. Various methods can be used to find the rules and estimate the quantities. The final choice is up to the particular problem and intended audience. The next section shows our CRL framework to achieve these goals.

### v3

@@ Abstract
Interpretability plays a crucial role in the application of statistical learning to estimate heterogeneous treatment effects (HTE) in complex diseases. In this study, we leverage a rule-based workflow, namely causal rule learning (CRL), to estimate and improve our understanding of HTE for atrial septal defect, addressing an overlooked question in the previous literature: what if an individual simultaneously belongs to multiple groups with different average treatment effects? The CRL process consists of three steps: rule discovery, which generates a set of causal rules with corresponding subgroup average treatment effects; rule selection, which identifies a subset of these rules to deconstruct individual-level treatment effects as a linear combination of subgroup-level effects; and rule analysis, which presents a detailed procedure for further analyzing each selected rule from multiple perspectives to identify the most promising rules for validation. Extensive simulation studies and real-world data analysis demonstrate that CRL outperforms other methods in providing interpretable estimates of HTE, especially when dealing with complex ground truth and sufficient sample sizes.
@@ 1 Introduction
In clinical research and practice, patients respond differently to specific treatment interventions, necessitating the exploration of heterogeneous treatment effects (HTE) to inform personalized treatment plans and optimize health outcomes. From a practitioner’s perspective, this exploration requires not only accurate estimation but also a deep understanding of HTE, which can be challenging to achieve in real-world clinical settings [crab2022benchmarking], especially for complex diseases such as atrial septal defect (ASD). As a common type of congenital heart disease, ASD exhibits diverse pathologies and patient heterogeneity [brida2021atrial]. With an accurate estimation of HTE and improved interpretation of HTE estimates, cardiologists can better understand, trust, and justify their treatment recommendations, thereby encouraging effective doctor-patient communication and collaboration [herlitz2017comparativism, crab2022benchmarking].
@@
So far, researchers have made great efforts to estimate HTE accurately using various methods, among which statistical learning methods are particularly popular due to their ability to handle high-dimensional datasets and uncover non-linear relationships between covariates [wager2018estimation, zhao2012estimating, seibold2016model, zhou2017residual]. However, most methods are typically designed for general estimation purposes, lacking specific adaptations for clinical applications. Besides, they seldom provide guidance on how to interpret what their model has learned from the data. Consequently, clinicians may struggle to understand these methods, let alone apply them in real-world practice.
@@
There has been an increasing amount of research addressing the interpretability issue in HTE estimation [dwivedi2020stable, hapfelmeier2018subgroup, subgrp2017huang, wang2021causal, chen2017general, foster2011subgroup]. The majority of this research focuses on identifying subgroups with enhanced (positive) treatment effects. However, the setup and methodology of these studies may be far from sufficient to reveal the complex nature of certain diseases. For complex disease treatment whose outcome is influenced by a lot of factors and interactions among these factors, we may find the situation shown in Figure 1: There are $M$ true subgroups (from subgroup 1 to subgroup $M$) that contribute to treatment effect heterogeneity. The population within each subgroup (e.g., , subgroup 1) shares the same treatment effect ($\tau_{1}=20$), i.e., the subgroup’s conditional average treatment effect (CATE) derived from the difference between the two treatment groups within the subgroup. The subgroup CATEs between groups vary in effect size (or magnitude) and sign ($\tau_{2}=15$ and $\tau_{3}=-8$). Besides, these subgroups are characterized with (not necessarily) different features, such as subgroup 1 by $chdhis\leq 1$ and $diabp\leq 58$ or subgroup 2 by $NYHA\leq 1$ and $BMI\leq 16.02$. Therefore, when we find the individual belongs to subgroups 1, 2, and $M$ ($r_{1}=1,r_{2}=1,r_{M}=1$) at the same time (multiple memberships), it raises the question of what the individual-level treatment effect (ITE) should be, given that the three subgroups differ in treatment effect. Considering the complex nature of certain diseases, we assume ITE as a linear combination of the above candidate group CATE in this study. Hence, the ITE in Figure 1 is $1/3\times 20+1/5\times 15+1/10\times(-7)=173/21$ where $\beta_{1}=1/3$, $\beta_{2}=1/5$, and $\beta_{M}=1/10$ are the linear coefficients in the combination which quantifies the contribution of each candidate subgroup to ITE.
@@
Compared with our perspective, the studies focus on subgroup identification usually end up with a limited number of subgroups for interpretability concerns [chen2023estimating, vegetabile2021distinction] (e.g., only subgroups 1 and 2 in Figure 1 are identified). Some approaches even define mutually disjoint subgroups (e.g., [wang2021causal]): One individual can only belong to one subgroup. As indicated in [vegetabile2021distinction], individual effects are not guaranteed to have even the same sign as CATE, let alone the same magnitude. Therefore, an ITE cannot be determined solely by belonging to a few specific groups, which relates to a few covariates. On the contrary, ITE may involve a broader range of subgroup memberships, and we need to consider how being in these subgroups collectively influences an individual, i.e., to explore the relationship between individual-level treatment effects and group-level treatment effects. In addition, since these studies emphasize a few subgroups, they are likely to ignore the estimation and understanding of ITE for the population outside the identified subgroups.
@@
With consideration of the above, this study introduces a novel method, Causal Rule Learning (CRL), to integrate both accurate estimation and an in-depth understanding of HTE for complex diseases. Guided by the Predictive, Descriptive, Relevant (PDR) framework of interpretability [yubinPDR2019] (see Appendix A for more information) and inspired by the RuleFit [friedman2008predictive] methodology, CRL caters to healthcare practitioners via an easily understood rule-based approach and an enlightening perspective on the relationship between ITE and CATE. Specifically, CRL incorporates three steps, i.e., rule discovery, rule selection, and rule analysis. The rule discovery step first generates a set of causal rules (characterized by different patient covariates) and their corresponding subgroup CATE estimates using causal forest [wager2018estimation]. Then, the rule selection step leverages a D-learning method [qi2018d] to filter out non-informative rules and use the informative ones to estimate ITE. Lastly, the rule analysis step evaluates the rules from three different granularity levels, namely overall analysis, significance analysis, and decomposition analysis. The whole workflow with details is shown in Figure 2.
@@
Figure 1: Connection between individual-level and subgroup-level treatment effects.
@@
CRL contributes to existing literature and healthcare practice by enriching the toolset for interpretable HTE estimation, particularly in scenarios involving complex disease treatments. In detail, we propose:
@@
    A comprehensively interpretable framework for HTE estimation. CRL presents a rule-based workflow that includes the discovery, selection, and analysis of rules. Rules are used for their ease in human understanding, usefulness in healthcare practice, and inclusion of potential knowledge. The discovery step seeks a large number of rules, which are meaningful for revealing the unknown aspects of HTE. The following selection step imposes sparsity on these rules, surfacing the most predictive and promising ones for ITE and further analysis. The analysis step provides an in-depth investigation of the rules identified, ranging from the overall performance of the rules as a whole on HTE-related tasks to how each component contributes to a specific rule in differentiating effect magnitudes, providing an in-depth interpretation of the subgroups identified.
@@
Together, the refined rules, their corresponding subgroup CATEs, and weights (coefficients in the linear combination) obtained after the whole workflow provide insights to HTE estimation and benefit relevant clinical practice, such as inspiring the stratification design of related RCTs and clinical decision support systems, as illustrated later in our real-world application of CRL.
The rest of this paper is organized as follows: In Section 2, we briefly summarize previous studies on HTE for clinical purposes. Section 3 delineates the proposed CRL workflow, including its basic data settings and assumptions, formulation, and workflow. Section 4 demonstrates the effectiveness of CRL on both simulated data and real-world data. Section 5 concludes our work and discusses its limitations and potential future research directions. For brevity in this paper, we will slightly misuse the following terms: i) HTE, CATE, and ITE; ii) satisfy a rule and belongs to a subgroup. These terms are used according to specific contexts and sometimes indiscriminately in broader situations.
Figure 2: Workflow of causal rule learning.
@@ 2 Related work
This section presents extant literature related to HTE estimation from various perspectives. The majority of studies focus on the estimation of CATE, namely the average treatment effect on specific subgroups. A small stream of research focuses on treatment recommendation, i.e., directly modeling the best individualized treatment rules or policy given personal characteristics to maximize the outcome of interest. Weighting methods are frequently used for this goal, such as [zhao2012estimating, HE2021107167, zhou2017residual, li2024penalized]. Though the latter is capable of providing actionable suggestions to clinical treatment decisions, they often fail to give explicit decision rationale or insights into HTE, which is really crucial for complex disease treatment. Therefore, we mainly discuss the studies on CATE estimation below.
CATE research generally has two focal points. One is developing general CATE estimators with favorable statistical properties using statistical methods such as single-index models [song2017semiparametric, FENG2022107554] or (causal) machine learning methods such as causal forest [athey2019generalized] and meta-learners [he2022x, polley2011super, van2011targeted]. Some studies involve complicated computations or even black-box models [johansson2016learning, yoon2018ganite, alaa2018limits] that are inherently unexplainable to improve estimation accuracy, failing to provide sufficient interpretation catering to clinical practitioners. The other focus emphasizes identifying and characterizing subgroups with statistically significant and practically meaningful heterogeneity in treatment response, such as [chen2015prim, su2009subgroup]. Tree-based methods such as [hapfelmeier2018subgroup, lipkovich2011subgroup, seibold2016model, su2009subgroup] have a natural advantage for this end due to their heuristic mechanism on population segmentation.
In particular, there is a growing field of CATE studies with rule ensemble models that follow quite different philosophies and workflows. As examples, [dwivedi2020stable] proposes a method called Stable Discovery of Interpretable Subgroups via Calibration to discover interpretable and stable rule sets from well-calibrated samples across 18 popular CATE estimators. Similarly, [bargaglistoffi2024causal] uses multiple CATE estimators as outcomes to generate tree ensembles for rule generation and then prunes the rules with certain penalties. They re-estimate CATE for the subpopulation characterized by the remaining rules. [wang2021causal] proposes Causal Rule Sets to identify subgroups with enhanced treatment effect through a generative Bayesian framework that assumes a prior for the actual rule set and a Bayesian logistic regression model to improve it. As mentioned earlier, all these studies presume a limited number of true subgroups, which may lead to the truth of complex disease being under-explored, especially for the population outside the identified subgroups.
@@ 3.1.1 Notation and assumptions
Suppose that we have observed a data set $\mathcal{D}=\{X_{i},A_{i},Y_{i}\}^{N}_{i=1}$ which comprises $N$ independent and identically distributed samples. This data set may come from an observational study or, ideally, a randomized controlled trial (RCT) setting. The proposed framework can be applied to data from both settings but requires the satisfaction of the three basic assumptions of the Neyman–Rubin potential outcome framework [rubin1980randomization, rosenbaum1983central], namely the Stable Unit Treatment Value Assumption, Unconfoundedness Assumption, and Overlap Assumption. Details of these assumptions are given in Appendix B.
3.1.2 ITE based on causal rules (subgroup CATEs)
In this study, we estimate ITE directly through a set of causal rules, by which we mean that the rules are identified from a causal analysis framework. As a natural way to represent human knowledge and a widely-used tool in clinical guidelines, they uncover variables’ potential interaction (synergistic effect) and define certain subgroups to help us understand the heterogeneity of the treatment effect. Most importantly, [friedman2008predictive] and [yubinPDR2019] show evidence that rule-based models have comparable and sometimes even better performance than black box methods like neural networks. Therefore, rules match our needs for both accuracy and interpretability. This paper formulates a rule as:
@@ 3.1.2 ITE based on causal rules (subgroup CATEs)
where $x_{j}$ denotes the $j$-th element of $X$, $s_{j}$ is a specified subset of all values of $x_{j}$, and $\mathop{1}(\cdot)$ is the indicator function yielding 1 when $x_{j}\in s_{j}$ and 0 otherwise. For example, $x_{1}>0$ and $x_{2}<0$ and $x_{3}>5$ is a rule representation consisting of three components. Each component (e.g., $x_{1}>0$) has a splitting variable ($x_{1}$) and a given cut-off value or threshold (0), which defines $s_{1}$. A rule $r(X)$ maps an input $X$ into $\{0,1\}$, representing the satisfaction of a rule with 1 and 0 otherwise. Note that a rule is usually represented only by the variables that help to split the whole population. For those variables that do not contribute to splitting the population, we assume $s_{j}=\mathbb{R}$.
@@
where $r_{m}(X)$, $\tau_{m}$, and $\beta_{m}$ denote the satisfaction of the $m$-th rule (0 or 1) given $X$, its corresponding subgroup CATE, and weight in the linear combination, respectively. Each subgroup CATE makes a different contribution to ITE, as reflected by the sign and magnitude of the weights $\beta$} and the subgroup CATEs. More significant absolute values of these quantities imply a greater contribution. It should be noted that (2) only models an additive relationship between the potential subgroups and does not account for more complex non-linear interactions between groups. We use this assumption mainly for clinical interpretability and simplicity for the demonstration of the CRL framework. For more complex scenarios, we can easily incorporate potential nonlinear interactions into (2) to capture more refined patterns of HTE. For instance, $\tau(X):=\sum_{m=1}^{M}\beta_{m}\tau_{m}r_{m}(X)+\sum_{i<j}\gamma_{i,j}f(\tau_{i}r_{i}(X),\tau_{j}r_{j}(X))$ where $\gamma_{i,j}$ is the coefficient for the between-group interaction term between group $i$ and group $j$.
Given above, the problem then boils down to the revelation of $\{r_{m}(X),\tau_{m},\beta_{m}\}^{M}_{m=1}$. Various methods can be used to find the rules and estimate the quantities. The final choice depends on the particular problem and intended audience. The next section shows our CRL framework to achieve these goals.